A golf ball with mass 45.90 g and diameter 42.60 mm is struck such that it moves with a speed of 54.15 m/s while rotating. The golf ball has a kinetic energy of 67.67 J. What is the rotational frequency of the golf ball (in rpm)?
1
Expert's answer
2020-04-28T09:46:44-0400

The total kinetic energy of the ball is the sum of the linear kinetic energy and angular kinetic energy:

"K_{tot} = K_{linear}+K_{angular}".

By definition:

"K_{linear} = \dfrac12mv^2; \space K_{angular} = \dfrac12I\omega^2" ,

where "m" is body mass, "v" - linear velocity, "I" - moment of inertia around the axis of rotation, "\omega" - angular rotational frequency.

The moment of inertia of the ball is "I = \dfrac{1}{10}md^2" , where "d" is the diameter of the ball.

Putting it all together in the first expression, obtain:

"K_{tot} =\dfrac12mv^2+ \dfrac12\dfrac{1}{10}md^2\omega^2" .

Express "\omega" and substitute numerical values:

"\omega^2 =\dfrac{ 20(K_{tot}- \dfrac12mv^2)}{md^2}=\dfrac{ 20(67.67- 0.5\cdot 0.0459\cdot 54.15^2)}{0.0459\cdot 0.0426^2} = 90157" .

Thus, "\omega = \sqrt{90157} \approx300" rad/s

The rotational frequency in rpm is:

"f = 60\cdot\dfrac{\omega}{2\pi} = 2864.8" rpm.


Answer. 2864.8 rpm.